A 130μF130μF capacitor is connected to a source that operates at 90Hz90Hz with a peak potential difference of 8V8V. Find:


d) the instantaneous power delivered to the capacitor at 4ms
Answer #1

The instantaneous power delivered to the capacitor,

P_{inst}(t) = \frac{1}{2}*2 \pi fCV^2 \cos (2 *2 \pi f t + \frac{\pi}{4}) = \pi fCV^2 \cos (4 \pi f t + \frac{\pi}{4})

P_{inst}(4 ms) = \pi *90*130*10^{-6} * 8^2 \cos (4 \pi *90*4*10^{-3} + \frac{\pi}{4}) = 1.32 W